OpenStudy (anonymous):

A hotel room cost $88 per night and you are getting 2 rooms for 6 nights for the group on your trip. What is the total hotel bill, without tax and fees? $1056 $902 $880 $1120

OpenStudy (anonymous):

i think is A

OpenStudy (anonymous):

88 x 2 = 176 176(6) = 1056 (A)

OpenStudy (anonymous):

you are right

OpenStudy (anonymous):

What was your speed if you drove 325 miles in 5 1/4 hours? 58 mph 62 mph 60 mph 65 mph

OpenStudy (jacob902):

(35*4)+(0.79*389) Multiply 35 by 4 to get 140. 140+(0.79*389) Multiply 0.79 by 389 to get 307.31. 140+307.31 Add 307.31 to 140 to get 447.31. 447.31 2(88*6) Multiply 88 by 6 to get 528. 2*528 Multiply 2 by 528 to get 1056 1056

OpenStudy (anonymous):

325/5.25 = about 62 (B)

OpenStudy (anonymous):

mileage Baltimore, MD Binghamton, NY Gettysburg, PA Philadelphia, PA Trenton, NJ Allentown, PA 177 134 123 66 78 Harrisburg, PA 88 184 39 114 133 Philadelphia, PA 102 184 143 35 Reading, PA 112 158 96 65 84 Scranton, PA 198 59 159 126 138 How far is it from Philadelphia, PA to Binghamton, NY? 184 mi 123 mi 158 mi 96 mi

OpenStudy (anonymous):

184 (A)

OpenStudy (anonymous):

just find the corresponding values on the chart

OpenStudy (anonymous):

Find the distance driven, if you drive at 60 mph for 3 1/2 hours. 17 mi 210 mi 180 mi 215 mi

OpenStudy (anonymous):

60 x 3.5 = 210 (B)

OpenStudy (anonymous):

If it cost $4.65 per day for a 5 day work week to ride the bus, and a weekly pass is $20, how much could you save by buying the pass? $3.00 $3.25 $3.50 $4.00

OpenStudy (jacob902):

D = 60 * 3.5 D = 210 miles

OpenStudy (anonymous):

4.65 x 5 = 23.25, if the weekly pass refers to a work week then (B) 23.25 - 20 = 3.25

OpenStudy (anonymous):

Farris Taxi Co. Fare First 1/4 mile $1.25 Each additional 1/4 mi $0.35 Each additional passenger $1.25 Find the total fare with 15% tip: 1 passenger, 5 1/2 miles. $8.60 $9.20 $9.89 $10.23

OpenStudy (jacob902):

4.16 times 5 is $23.25 so you would save $3.25 per week

OpenStudy (jacob902):

The passenger will pay 1.25 for the first .25 miles of the trip. This leaves 5.25 miles to be paid at a rate of $.35 per quarter mile (5.25 x 4 = 21 quarter miles) Fare =1.25 + .35 [(4) (5.5 - .25)] Fare = 8.60 Cost (including 15% tip): Fare x 1.15 = 8.60 x 1.15 = 9.89 Total cost is $9.89.

OpenStudy (anonymous):

(C)

OpenStudy (anonymous):

is MST Mountain?

OpenStudy (texaschic101):

I think so

OpenStudy (anonymous):

(C)

OpenStudy (anonymous):

2 hrs 50 mins

OpenStudy (anonymous):

QUESTION 52 What is the price per oz to the nearest cent of the least expensive: 15 oz for $4.56 or 25 oz for $5.98? 0.28 0.25 0.3 0.24

OpenStudy (anonymous):

ill get in one sec

OpenStudy (texaschic101):

4.56/15......do this and it will give you price per oz 5.98/20......= price per oz now, what is the least one

OpenStudy (anonymous):

D

OpenStudy (anonymous):

.239 is the second one

OpenStudy (anonymous):

.304 is the first

OpenStudy (anonymous):

3000 medals wtf

OpenStudy (anonymous):

so number 52 is D??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

QUESTION 51 You have a $20 bill. You and your friend ordered two cheeseburgers for $3.86 each, two fries for $1.68 each, and two colas for $2.15 each. Including 6% sales tax and 15% tip, what is the change you will receive from $20? $4.62 $3.52 $2.39 $1.39

OpenStudy (anonymous):

ugh lol

OpenStudy (anonymous):

is tap added to total cost plus sales tax?

OpenStudy (anonymous):

tip*

OpenStudy (anonymous):

D

OpenStudy (anonymous):

51 is d?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

QUESTION 50 Fleece vests regularly priced at $15.99 are on sale for $13.99. If the sales tax is 5%, how much can you save by buying the vests on sale? $2.00 $2.10 $2.20 $2.30

OpenStudy (anonymous):

B

OpenStudy (anonymous):

QUESTION 49 You bought a coat at $78.50 and two shirts at $21.99 each. If the sales tax rate is 7%, how much is the total purchase? $122.36 $125.65 $127.35 $131.05

OpenStudy (anonymous):

D

OpenStudy (anonymous):

so do i have 48 questions left?

OpenStudy (anonymous):

QUESTION 48 Find the percent cost of the total spent on each: Equipment $36, Fees $158, Transportation $59 14%, 62%, 23% 15%, 60%, 35% 10%, 70%, 20% 11%, 62%, 27%

OpenStudy (anonymous):

A

OpenStudy (anonymous):

they sum up to 99 cuz technically 62 percent should be 63 i think but A is right

OpenStudy (anonymous):

QUESTION 47 Mary bought these art supplies: 4 paint brushes @ $1.29 each, 6 dowel rods @ $1.49 each, a jar of paste @ $1.25, and 2 canvases @ $6.00 each. How much change did she get from $40? $12.65 $10.45 $8.25 $9.35

OpenStudy (anonymous):

im getting off sooner or later XD

OpenStudy (anonymous):

ohh that bad =(

OpenStudy (anonymous):

hahahaha dont u know how to do these XD

OpenStudy (anonymous):

im hispanic don't know english

OpenStudy (anonymous):

A

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

QUESTION 46 The total cost of dinner at a restaurant for Shelly and her 4 friends was $138.60. If each person paid the same amount, how much did each person pay? $34.65 $27.72 $31.91 $32.45

OpenStudy (anonymous):

B

OpenStudy (anonymous):

46 is B??

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

QUESTION 45 Alton had $18.00 to spend on paints, which sold for $1.14 each. How many jars of paint can he buy? 18 17 16 15

OpenStudy (anonymous):

ima go after 40

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so if u see a harder one give me that plz if u actually cant do it

OpenStudy (anonymous):

D

OpenStudy (anonymous):

QUESTION 3 Find the hourly wage: $70 for 9 hours. $71 $7.78 $17.78 $10.12 2 points QUESTION 4 These three jobs were in the paper today: Frank's Furters $5.67 /hr. Harriet's Hair Styles $80 for 15 hours weekly. Babysitting $250 a month for 10 hours after school a week. (Assume 4 weeks is a month.) Which job has the highest hourly pay? Frank's Furters Harriet's Hair Styles Babysitting Two of them are tied. 2 points QUESTION 5 You just worked a week for Jan's Jewels. Your pay before deductions was $101.46 and you worked 19 hours. How much did you make per hour? $5.16 $4.90 $5.34 $6.88 2 points QUESTION 6 You have just accepted a job at Wonders Day Camp for the summer. The camp will pay you $37.25 each day for 10 weeks. If you work 5 days each week, how much money will you make this summer? $1,490.50 $1,298.75 $1,192 $1,862.50 2 points QUESTION 7 Write as a percent: 5/8 0.625% 6.25% 62.5% 625% 2 points QUESTION 8 Margaret said that 1/5 of her class was absent today. What percent of the class was absent? 10% 15% 20% 30% 2 points QUESTION 9 Write as a decimal: 46% 0.46 4.6 460 0.046 2 points QUESTION 10 Write as a percent: 0.564 5.64% 56.4% 0.564% 564% 2 points QUESTION 11 Write as a fraction in lowest terms: 40% 40/100 2/5 4/10 8/20 2 points QUESTION 12 Arthur is treating his friends to a meal at Chez Rich. The meals added up to $136.45. If he tips the waiter 15%, what is the amount of the tip? $20.46 $20.47 $13.36 $13.37 2 points QUESTION 13 13% of $67.99 = $8.84 $8.16 $7.53 $6.58 2 points QUESTION 14 Evaluate: 16 - 2 + 3 - 4 1 17 9 13 2 points QUESTION 15 Compute: (4+9) - 3(6)/2 + 8 12 38 11.2 4 2 points QUESTION 16 Compute: 2(8) + 5(6) = 126 40 46 24 2 points QUESTION 17 Compute the total pay: Hours worked: 6 Hourly wage: $3.95 Plus Tips: $21.65 $33.00 $65.50 $45.35 $33.85 2 points QUESTION 18 Find the percent of tip: Cost of meal $18.50, Tip $2.59 13% 14% 15% 16% 2 points QUESTION 19 If 3/4 of the senior class voted to take a field trip and the class had 535 members, how many voted for the trip? 134 255 401 562 2 points QUESTION 20 Fortuna worked 49.5 hours this week at $8.18 an hour. What was her gross pay? $404.91 $396.73 $289.54 $283.88 2 points QUESTION 21 Time in Time out 7:05 3:58 7:30 5:01 8:01 5:32 7:22 4:46 6:55 4:31 Margie has these times recorded on her time sheet for the week. How many hours did she work? 47 37 45 35 2 points QUESTION 22 Write as a mathematical expression. The product of 5 and x. 5+x 5 - x 5x 5/x 2 points QUESTION 23 Write an algebraic expression for the word phrase. 8 more than the product of 6 and n. 6(8 + n) 8(6 + n) 8 + 6 + n 8 + 6n 2 points QUESTION 24 Write as an algebraic expression: 22% of a. 0.22a 0.22 + a 0.22 / a 0.22 - a 2 points QUESTION 25 Evaluate ab + c for a= 2, b = 5, c = 4. 18 22 14 20 2 points QUESTION 26 Solve: x - 2.8 = 1.6 -1.2 4.4 0.57 3.4 2 points QUESTION 27 Art bought 4 boxes of pens. After Art gave his friend Ben 8 pens, he has a total of 112 pens. How many pens are in a box? Which of the following is the best equation for the problem? 4X - 8 = 112 4X + 8 = 112 X /4 - 8 = 112 112 - X/4 = 8 2 points QUESTION 28 Your regular pay of $67, plus tips, came up to $182.50. How much was the tips? $114.25 $115.50 $67 $125.50 2 points QUESTION 29 Solve: 3x + 18 - 7 = 41 10 27 33 52 2 points QUESTION 30 Solve: y ÷8 + 19 = 30 88 392 1.375 43.5 2 points QUESTION 31 Solve: 14 = 0.1x 1.4 14 140 1400 2 points QUESTION 32 If Job A pays $459 semimonthly and Job B pays $449 biweekly, which will pay you more? Job A Job B They both pay the same. 2 points QUESTION 33 If you earn a commission of $150.00 per course that you sell on the internet, and you have sold 112 courses, what is your commission? $15,000 $18,150 $13,550 $16,800 2 points QUESTION 34 Find the Net Pay: $678 gross pay, Federal Withholding $132, FICA $51.87, State $43.15, Health $13.25. $437.73 $653.67 $701.73 $678 2 points QUESTION 35 Deductions this Week: Federal Tax $52.02, FICA $43.02, State $5.35. Find the annual Federal Tax deduction. $2,705.04 $1,717.04 $4,422.08 $277.16 2 points QUESTION 36 Find the Take-Home percent of the Gross Pay: Net Pay $453.10, Gross Pay$734. 50% 16% 162% 62% 2 points QUESTION 37 Your health insurance premium is $4211 annually. If your employer pays 80% and you are paid monthly, how much is deducted from each monthly check? $16.20 $70.18 $842.20 $900.00 2 points QUESTION 38 Annual Premium per $1000 Age Smoker Age Non-Smoker 10 year 20 year 20 25 2.55 3.60 25 30 3.12 5.74 30 35 3.80 7.85 35 40 4.25 9.36 Some insurance companies charge more if you smoke cigarettes. What is the difference in the annual life insurance premiums paid by a 35 year old smoker and a 35 year old non-smoker, if they both take out a $100,000 policy for 20 y term? $151 $259 $765 9.35 2 points QUESTION 39 How many more SUV's were leased than 2 Door Sedans? Imported Asset 40 45 50 55 2 points QUESTION 40 Which year had the highest month of car sales? Imported Asset 2004 2005 neither 2 points QUESTION 41 Find the AGI and taxable income: Gross Income $23,670 Adjustments 0 1 Exemption $8,200 Deduction 0 $23,670 and $8,200 $23,670 and $15,470 $15,470 and $8,200 $31,870 and $23,670 2 points QUESTION 42 Imported Asset Carol is single and has an adjusted gross income of $32,785. She claims one exemption of $8,200. What is her tax due? $3,314 $2,949 $3,321 $2,958 2 points QUESTION 43 You are going to movies Friday night (not a matinee). Which of the theaters will cost the least, if you intend to have soda and popcorn? Theater A Theater B Theater C Ticket $6.25 $7.25 $5.25 Soda $2.00 $2.00 $2.50 Popcorn $2.50 $2.00 $2.50 Theater A Theater B Theater C tie between two of them 2 points QUESTION 44 You and 3 friends are going to Great Mountain Park. The park has 3 water rides. The Coaster, Slide, and Drop cost $5.50, $2.00, and $6.50, respectively. Your group plans to ride each water ride once. How much change will you get back from a $100 bill? $14 $22 $50 $44 2 points QUESTION 45 Alton had $18.00 to spend on paints, which sold for $1.14 each. How many jars of paint can he buy? 18 17 16 15

OpenStudy (anonymous):

QUESTION 44 You and 3 friends are going to Great Mountain Park. The park has 3 water rides. The Coaster, Slide, and Drop cost $5.50, $2.00, and $6.50, respectively. Your group plans to ride each water ride once. How much change will you get back from a $100 bill? $14 $22 $50 $44

OpenStudy (anonymous):

D

OpenStudy (anonymous):

QUESTION 43 You are going to movies Friday night (not a matinee). Which of the theaters will cost the least, if you intend to have soda and popcorn? Theater A Theater B Theater C Ticket $6.25 $7.25 $5.25 Soda $2.00 $2.00 $2.50 Popcorn $2.50 $2.00 $2.50 Theater A Theater B Theater C tie between two of them

OpenStudy (anonymous):

wtf

OpenStudy (anonymous):

ok ill do this big chunk that starts with question 3 then im gonna go

OpenStudy (anonymous):

it will take a bit

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

your the best hazer

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

in regards to question 7, 5/8 out of 1 or 100? im going to assume 100 for now tell me otherwise

OpenStudy (anonymous):

same with number 10

OpenStudy (anonymous):

oh nevermind i think its one becuz 64% is .64

OpenStudy (anonymous):

ask someone else number 38 please, i dont know it XD

OpenStudy (anonymous):

also with 39 not enough info is provided

OpenStudy (anonymous):

q3 = B q4 = c q5 = c q6 = d q7 = c q8 = c q9 = a q10 = b q11 = b q12 = b q13 = a q14 = d q15 = a q16 = c q17 = c q18 = b q19 = c q20 = a q21 = a q22 = c q23 = d q24 = a q25 = c q26 = b q27 = a q28 = b q29 = a q30 = a q31 = c q32 = a q33 = d q34 = a q35 = a q36 = d q37 = b q38 = ? q39 = ? q40 = ? q41 = ? q42 = ? q43 = c q44 = d q45 = d

OpenStudy (anonymous):

Hope you still want them or even notice, if not i would be VERY MAD. so look at them XD